# Expert System for Bitcoin Forecasting: Integrating Global Liquidity via TimeXer Transformers

### v1

@@ 4.3 Training Setup
No manual normalization or scaling was applied during preprocessing. Instead, standardization was handled internally by each model through built-in normalization mechanisms such as Reversible Instance Normalization (RevIN) (Kim2022) used in PatchTST. This approach ensures that normalization statistics are computed independently for each sample, thereby eliminating potential information leakage across time.
@@
4.4 Implementation Using the NeuralForecast Library.
All deep learning models were implemented using the NeuralForecast library, which provides a state-of-the-art framework for reproducing and training deep learning forecasting models (Olivares2022). The NeuralForecast framework was used to train and evaluate all the models (TimeXer, PatchTST, LSTM, and N-BEATS).
The library has native support for future exogenous variables, thus we could easily integrate lagged Global Liquidity features into the TimeXer-Exog model. Model training was conducted using horizon-specific fitting with early stopping based on validation loss Rolling-origin forecasting was employed to assess true out-of-sample performance. This standardized implementation ensured fair comparison across all the models and minimized confounding effects arising from differences in training or evaluation procedures.
@@ 4.5 Evaluation Metrics and Forecast Horizons
Performance was evaluated on the held-out out-of-sample test window using Mean Squared Error (MSE) as the evaluation metric. For numerical clarity and comparability across horizons, all the MSE values are scaled by a factor of $10^{7}$, with lower values indicating better predictions.
@@ 5.1 Quantitative Performance
TimeXer-Exog begin to dominate after the one month mark. From 28 days onward, TimeXer-Exog has the lowest MSE and has great stability. There is less drifts and there seems to be a clearer sense of direction. Univariate models such as LSTM and NBEATS, performed poorly as the forecast horizon increased. Without outside information, their predictions either become flat or drift aimlessly, therefore showing the limits of using only price data to forecast.
@@ 5.2 Qualitative Forecast Analysis
Figure 3: Forecast comparison at horizon $H=7$.
@@
Figure 4: Forecast comparison at horizon $H=28$.
@@
Figure 5: Forecast comparison at horizon $H=63$.
@@ 5.3 Statistical Significance and Robustness Analysis
Mean squared error is an easy way to rank models from best to worst. Still, it only scratches the surface. Prices jump for reasons that have little to do with forecasting skill, and a tiny edge in error can just as easily be luck as insight. That uncertainty tends to get glossed over when models are ranked mechanically.
This is where the Model Confidence Set becomes useful. Table 2 reports the corresponding p-values across different forecast horizons, which forces a pause before declaring winners. Within this setup, a p-value of 1.000 means a model survives every comparison and stays in the set of top performers. Lower values tell a less flattering story. Once the threshold drops below 0.10, the supposed edge starts to look fragile, the kind that might disappear if the sample were reshuffled or extended by a few months. At that point, exclusion is hard to avoid.
Seen this way, the procedure does not magically identify the “true” best model. Rather, it reins in overconfidence. It helps distinguish improvements that persist under repeated scrutiny from those that shine briefly, perhaps because a particular market phase happened to line up in their favor. That distinction matters, especially when forecasting models are expected to hold up beyond the specific slice of data that first made them look good.
Table 2: Model Confidence Set (MCS) P-Values across Forecasting Horizons. Bold values indicate inclusion in the 90% confidence set ($\alpha=0.10$). Models with $p<0.10$ are statistically significantly inferior.
@@
Long-Term Superiority ($h\in[56,70]$): Things change as the horizon increases. For forecasts at $h=63$ and $h=70$, the results in Table 2 are unambiguous. TimeXer-Exog is the only one in the confidence set, with a p-value of 1.000. All other models, including the base TimeXer and PatchTST, fall out well before the 10 percent threshold. At this point, the advantage of incorporating exogenous information is clearly evident.
Discussion of Results: This separation at longer horizons speaks directly to the role of global M2 liquidity. As forecasts horizons increase, the usefulness of past price movements fades. Autoregressive signals weaken, and univariate models begin to drift, often collapsing toward their own long-run averages or accumulating error along the way. The TimeXer-Exog model, drawing on a 12-week lagged liquidity signal, it retains access to information that is not present in recent prices. Because this signal leads the market rather than reacting to it, it offers a more stable reference point when price-based cues run thin. The MCS results suggest that the improvement in long-horizon accuracy is structural rather than accidental. In other words, the gains appear to come from how macroeconomic information is built into the model, not from statistical luck.
@@
All data used here are publicly accessible. The Bitcoin prices, along with the processed global liquidity series, can be found in the gihub repository at https://github.com/sravankarthik/bitcoin-timeXer/blob/main/bitcoin.csv . The same repository also contains the full notebook needed to replicate the analysis, including the steps used to construct the global liquidity measure, introduce lagged features, and train all the models. The TradingView Pine Script for bitcoin and global liquidity is included as well https://github.com/sravankarthik/bitcoin-timeXer . With these materials, the experiments reported in this paper should be reproducible in full, or at least close enough.
6 Conclusion
The results point to a fairly clear takeaway. Using global M2 liquidity as exogeneous variable in TimeXer helps turn what might otherwise be a prediction model into something closer to an expert system for digital asset management. Over longer horizons, up to about 70 days, anchoring forecasts to a macroeconomic signal appears to matter a great deal. The observed reduction in forecast error, close to 89 percent relative to standard benchmarks, suggests that the model is not simply fitting price patterns more tightly, but is aligning its expectations with broader monetary conditions.
These results are beyond just academic accuracy metrics. This framework could mitigate risk in risk management systems or exist in algorithmic trading pipelines. Using global liquidity we can prevent the model from overreacting to short-lived price shocks, the sort of sudden moves that often dominate crypto markets for a short time and then fade.
Long-horizon forecasting in crypto markets or any market seems unlikely to work if it relies only on past prices. Past prices are important, but without explicit links to macroeconomic drivers, forecasts tend to drift. Attention-based architectures that incorporate exogenous liquidity signals offer better way to connect predictions to underlying economic forces, which in turn makes them more resilient in highly volatile markets.

### v2

@@ 4.3 Training Setup
No manual normalization or scaling was applied during preprocessing. Instead, standardization was handled internally by each model through built-in normalization mechanisms such as Reversible Instance Normalization (RevIN) (Kim et al., 2022) used in PatchTST. This approach ensures that normalization statistics are computed independently for each sample, thereby eliminating potential information leakage across time.
@@
4.4 Hyperparameter Optimization
To confirm that the model configurations were not chosen arbitrarily, a systematic hyperparameter search using Optuna was performed, an automated optimization framework employing the Tree-structured Parzen Estimator (TPE) algorithm (Akiba et al., 2019). The TPE algorithm was selected for its efficiency in sampling high-dimensional search spaces as compared to traditional grid or random search methods.
Minimizing the MSE on the validation set was the primary optimization objective (September 24, 2024, to January 21, 2025). For the Transformer-based models (TimeXer and PatchTST), the search space was defined to explore various model depths, attention capacities, and temporal receptive fields. The specific ranges and discrete values used during the optimization process are listed in Table 2.
Table 2: Hyperparameter Search Space for Transformer-based Models.
Hyperparameter Search Space / Range
Input Size (Lookback) $\{128,256,512\}$
Patch Length $\{32,64,96,128\}$
Stride $\{8,16,32\}$
Hidden Size ($d_{model}$) $\{64,128,256\}$
Encoding Layers $\{4,8,12,16\}$
Attention Heads $\{4,8\}$
Dropout $[0.2,0.4]$ (Continuous)
Learning Rate $10^{-6},10^{-4}]$ (Log-uniform)
The optimization trials revealed that deeper architectures with larger patch sizes yielded the most stable long-term forecasts for Bitcoin. Specifically, a patch length of 96 with a stride of 8 provided the optimal balance between capturing local volatility and maintaining computational efficiency. Table 3 presents the final optimal configurations selected for each model based on the best validation performance.
Table 3: Final Optimized Hyperparameters for All Models.
Hyperparameter TimeXer-Exog TimeXer PatchTST LSTM N-BEATS
Input Size (Lookback) 256 256 256 256 256
Patch Length 96 96 96 - -
Stride 8 8 8 - -
Hidden Size ($d_{model}$) 128 128 128 256 (Enc) / 128 (Dec) 512 (MLP)
Layers 16 16 6 3 (Enc) / 2 (Dec) 3 Blocks
Attention Heads 8 8 8 - -
Dropout 0.30 0.25 0.20 - 0.30 0.15 0.0
FFN Dimension ($d_{ff}$) 512 512 512 - -
A consistent training configuration was employed specifically to ensure stable convergence across all architectures. All models were optimized using the Adam optimizer (Kingma and Ba, 2014), which is well-suited for non-stationary financial data due to its adaptive learning rate capabilities. We utilized a fixed batch size of 128, chosen to balance memory efficiency with gradient stability. Considering the high volatility of the Bitcoin price series, this batch size provided sufficiently robust gradient estimates to prevent erratic weight updates while maintaining high computational throughput on the GPU. Standard scaling (z-score normalization) was applied to inputs to further aid the optimizer in navigating the loss landscape.
4.5 Implementation Using the NeuralForecast Library
All deep learning models were implemented using the NeuralForecast library, which provides a state-of-the-art framework for reproducing and training deep learning forecasting models (Olivares et al., 2022). The NeuralForecast framework was used to train and evaluate all the models (TimeXer, PatchTST, LSTM, and N-BEATS).
The library has native support for future exogenous variables, thus we could easily integrate lagged Global Liquidity features into the TimeXer-Exog model. Model training was conducted using horizon-specific fitting with early stopping based on validation loss. Rolling-origin forecasting was employed to assess true out-of-sample performance. This standardized implementation ensured fair comparison across all the models and minimized confounding effects arising from differences in training or evaluation procedures.
@@ 4.6 Evaluation Metrics and Forecast Horizons
Performance was evaluated on the held-out out-of-sample test window using MSE as the evaluation metric. For numerical clarity and comparability across horizons, all the MSE values are scaled by a factor of $10^{7}$, with lower values indicating better predictions.
@@ 5.1 Quantitative Performance
TimeXer-Exog begins to dominate after the one month mark. From 28 days onward, TimeXer-Exog has the lowest MSE and has great stability. There are less drifts and there seems to be a clearer sense of direction. Univariate models such as LSTM and NBEATS, performed poorly as the forecast horizon increased. Without outside information, their predictions either become flat or drift aimlessly, therefore showing the limits of using only price data to forecast.
@@ 5.2 Qualitative Forecast Analysis
Figure 3: Forecast comparison at horizon $H=7$ days.
@@
Figure 4: Forecast comparison at horizon $H=28$ days.
@@
Figure 5: Forecast comparison at horizon $H=63$ days.
5.3 Interpretability Analysis: Visualizing Liquidity Transmission
To validate whether the performance gains at long-horizons stem from global liquidity signals, we inspected the cross-attention weights from the model’s first layer (Layer 0), averaged across all attention heads. Figure 6 visualizes which specific lags of Global Liquidity the model prioritizes when generating forecasts for different horizons. While attention patterns evolve across layers, we find that Layer 0 offers the clearest and most stable representation of liquidity–price transmission. Visual inspection of deeper layers revealed qualitatively similar but increasingly diffuse patterns; hence, we report Layer 0 for interpretability and clarity.
(a) Horizon $H=7$ (Short)
(b) Horizon $H=21$ (Medium)
(c) Horizon $H=63$ (Long)
Figure 6: Evolution of Cross-Attention Patterns (Layer 0). (a) At $H=7$, attention is focused on immediate indicators (global and global_lag_1). (b) At $H=21$, a distinct anchor emerges at global_lag_49 (7 weeks), suggesting a learned transmission delay. (c) At $H=63$, the focus shifts to deep lags (global_lag_84 to global_lag_91), aligning with the quarterly liquidity cycle.
The analysis reveals three distinct behavioral regimes:
1. Short-Term Reactivity ($H=7$): At the shortest horizon (Figure 6a), the model allocates the greatest attention to the most recent data points, particularly the target variable itself (Bitcoin_price) and immediate liquidity measures (global, global_lag_1). This confirms that for weekly forecasts, the model operates primarily as a momentum trader, relying on current market conditions rather than delayed macroeconomic signals.
2. Medium-Term Anchoring ($H=21$): A structural shift is observed at the 3-week horizon. As shown in Figure 6(b), the attention mechanism shifts away from current data and forms a distinct peak at global_lag_49. This indicates that for 21-day predictions, the model autonomously learned to ”look back” approximately 7 weeks. This 49-day lag likely captures the intermediate transmission period as liquidity changes filter through traditional financial systems before impacting risk assets.
3. Long-Term Alignment ($H=63$): At the longest horizons (Figure 6c), the attention weights shift to the deepest part of the context window, showing sustained focus on global_lag_84 and global_lag_91. This corresponds almost exactly to a 12-week (3-month) cycle. The model’s spontaneous reliance on these deep lags provides empirical support for the hypothesis that Global Liquidity acts as a leading indicator with a quarterly transmission lag, a relationship that simple univariate models fail to capture.
@@ 5.4 Statistical Significance and Robustness Analysis
Mean squared error is a simpler way to rank models from best to worst. However, it only scratches the surface. Prices jump for reasons that have little to do with forecasting skill, and a tiny edge in error can just as easily be luck as insight. That uncertainty tends to get glossed over when models are ranked mechanically.
This is where the MCS becomes useful. Table 5 reports the corresponding p-values across different forecast horizons, which forces a pause before declaring winners. Within this setup, a p-value of 1.000 means a model survives every comparison and stays in the set of top performers. Lower values tell a less flattering story. Once the threshold drops below 0.10, the supposed edge starts to look fragile, the kind that might disappear if the sample were reshuffled or extended by a few months. At that point, exclusion is hard to avoid.
From this perspective, the procedure does not magically identify the “true” best model. Rather, it reins in overconfidence. It helps distinguish improvements that persist under repeated scrutiny from those that shine briefly, perhaps because a particular market phase happened to line up in their favor. That distinction matters, especially when forecasting models are expected to hold up beyond the specific slice of data that first made them look good.
Table 5: MCS P-Values across Forecasting Horizons. Bold values indicate inclusion in the 90% confidence set ($\alpha=0.10$). Models with $p<0.10$ are statistically significantly inferior, meaning they are unlikely to be the best model.
@@
Long-Term Superiority ($h\in[56,70]$): Things change as the horizon increases. For forecasts at $h=63$ and $h=70$, the results in Table 5 are unambiguous. TimeXer-Exog is the only one in the confidence set, with a p-value of 1.000. All other models, including the base TimeXer and PatchTST, fall out well before the 10 percent threshold. This strongly suggests that incorporating exogenous information is very advantageous.
Discussion of Results: This separation at longer horizons points directly to the role of global M2 liquidity. The relevance of past price movements fades with an increase in forecast horizons. As autoregressive signals begin to weaken, univariate models tend to lose stability, drifting toward long-run averages or accumulating forecast errors over time. Drawing on a 12-week lagged liquidity signal, the TimeXer-Exog model retains access to information that is not present in recent prices. Rather than reacting dramatically to the movement of the market, this signal leads it, offering a more stable reference point when price-based cues weaken. The MCS results suggest that the improvement in long-horizon accuracy is structural rather than accidental. In other words, the gains appear to come from how macroeconomic information is built into the model, not from statistical luck.
6 Conclusion
The results demonstrate that integrating Global M2 Liquidity into the TimeXer architecture evolves the model from a passive statistical tool into a robust expert system for digital asset management. Over forecast horizons extending to 70 days, the system’s ability to anchor predictions to macroeconomic signals yields a reduction in forecast error of over 89% relative to univariate benchmarks. This performance gap confirms that the model successfully mimics expert intuition: it learns to prioritize slow-moving, high-impact monetary signals over short-term price noise when projecting long-term trends.
This framework serves as a critical decision-support tool for risk management and algorithmic trading. By explicating the relationship between global liquidity and asset prices through attention mechanisms, the system offers interpretability that purely technical models lack. It provides risk managers with “early warning” capabilities, preventing the model from overreacting to idiosyncratic price shocks while ensuring alignment with broader economic cycles. This capability allows for the construction of trading pipelines that reflect how sophisticated financial decisions are actually made—with one eye on market microstructure and the other on the global macroeconomic landscape.
Past prices are important, but without explicit links to macroeconomic drivers, forecasts tend to drift. Long-horizon forecasting in crypto markets or any market seems unlikely to work if it relies only on past prices. Attention-based architectures that incorporate exogenous liquidity signals offer a better way to connect predictions to underlying economic forces, which in turn makes them more resilient in highly volatile markets.
@@ 6 Conclusion
All data used here are publicly accessible. The Bitcoin prices, along with the processed global liquidity series, can be found in the GitHub repository at https://github.com/sravankarthik/bitcoin-timeXer/blob/main/bitcoin.csv . The same repository also contains the full notebook needed to replicate the analysis, including the steps used to construct the global liquidity measure, introduce lagged features, and train all the models. The TradingView Pine Script for bitcoin and global liquidity is included as well https://github.com/sravankarthik/bitcoin-timeXer . With these materials, the experiments reported in this paper should be reproducible in full, or at least close enough.